# Issue2Test: Generating Reproducing Test Cases from Issue Reports

### v1

@@ Abstract.
Automated tools for solving GitHub issues are receiving significant attention by both researchers and practitioners, e.g., in the form of foundation models and LLM-based agents prompted with issues. A crucial step toward successfully solving an issue is creating a test case that accurately reproduces the issue. Such a test case can guide the search for an appropriate patch and help validate whether the patch matches the issue’s intent. However, existing techniques for issue reproduction show only moderate success. This paper presents Issue2Test, an LLM-based technique for automatically generating a reproducing test case for a given issue report. Unlike automated regression test generators, which aim at creating passing tests, our approach aims at a test that fails, and that fails specifically for the reason described in the issue. To this end, Issue2Test performs three steps: (1) understand the issue and gather context (e.g., related files and project-specific guidelines) relevant for reproducing it; (2) generate a candidate test case; and (3) iteratively refine the test case based on compilation and runtime feedback until it fails and the failure aligns with the problem described in the issue. We evaluate Issue2Test on the SWT-bench-lite dataset, where it successfuly reproduces 30.4% of the issues, achieving a 40.1% relative improvement over the best existing technique. Our evaluation also shows that Issue2Test reproduces 28 issues that seven prior techniques fail to address, contributing a total of 68.3% of all issues reproduced by any tool. We envision our approach to contribute to enhancing the overall progress in the important task of automatically solving GitHub issues.
@@ 1. Introduction
Issue reports, e.g., on GitHub, are commonly used to describe bugs, missing features, and other ways to improve a software project. Because addressing issues takes significant developer effort, recent work invests heavily in automated issue solving. Motivated by benchmarks of issues in popular open-source projects, such as SWE-bench (Jimenez et al., 2024), foundation models and Large Language Model (LLM)-based software engineering agents compete in their ability to successfully solve issues (Zhang et al., 2024; Roychoudhury et al., 2025; Rondon et al., 2025; Yang et al., 2024; Wang et al., 2024b; Chen et al., 2024; Tao et al., 2025; Xia et al., 2024). In addition to approaches that directly work on issues, automated program repair (Le Goues et al., 2019) tries to address bugs that manifest through a failing test. Various techniques have been explored for automated bug repair (Fan et al., 2023; Xia and Zhang, 2022; Bouzenia et al., 2024).
A crucial prerequisite for successfully solving an issue or repairing a bug is a test case that reproduces the problem. Unfortunately, issue reports rarely include executable tests that reproduce the issue, making it challenging for developers and automated techniques to validate potential patches. Specifically in open-source projects, contributors often lack the expertise and familiarity to construct meaningful test cases alongside their reports. Most work on automated program repair repair assumes reproducin test cases to be given. In practice, however, developers often write test cases only after a problem was fixed (Cheng et al., 2025). Automated issue solving techniques sometimes include a step for reproducing the issue (Ruan et al., 2024), but do not explicitly focus on this important part.
One approach for obtaining an issue-reproducing test case could be automated test generation. Both traditional approaches (Fraser and Arcuri, 2011; Pacheco et al., 2007) and neural methods (Watson et al., 2020; Dinella et al., 2022) attempt to reduce the time spent on manually writing tests. However, earlier techniques often fail to generate human-readable and maintainable test cases that align with project conventions and effectively validate intended program behavior (Daka et al., 2017). The emergence of LLMs has significantly enhanced automated test generation (Ziegler et al., 2022; Vaithilingam et al., 2022; Moradi Dakhel et al., 2023). LLMs can synthesize tests with human-like readability and writing style, reflecting patterns learned from extensive training on large code repositories (Ziegler et al., 2022). This capability has prompted the use of LLMs in test code generation (Bareiß et al., 2022; Nashid et al., 2023; Lemieux et al., 2023; Siddiq et al., 2023; Schäfer et al., 2023b; Xie et al., 2023; Wang et al., 2024a). While useful for creating regression tests, none of these approaches addresses the problem of creating tests that reproduce a specific issue.
@@
We evaluate Issue2Test on all 276 instances of SWT-bench-lite, a subset of SWE-bench that has been extensively used for evaluating prior work (Mündler et al., 2024; Ahmed et al., 2024). The results show that our approach: i) outperforms all the existing baselines by generating tests for 84 instances (24 higher than the best baseline), ii) successfully generates tests for 28 unique issues missed by any prior work and contributes 68.3% of the union of reproduced issues by any technique, iii) imposes moderate costs of 5.21 cents per issue, which is orders of magnitude cheaper than what a human software developer would charge for performing the same task.
@@ 2.1. Running Example
Title Cannot override get_FOO_display() in Django 2.2+ Description I cannot override the get_FIELD_display function on models since version 2.2. It works in version 2.1. Example: 1class FooBar(models.Model): 2 foo_bar = models.CharField(_("foo"), choices=[(1, ’foo’), (2, ’bar’)]) 3 4 def __str__(self): 5 return self.get_foo_bar_display() # This returns ’foo’ or ’bar’ in 2.2, but ’something’ in 2.1 6 7 def get_foo_bar_display(self): 8 return "something" What I expect is that I should be able to override this function.
@@
Figure 2 shows a test case produced by Issue2Test for the issue in Figure 1, which successfully reproduces the described issue. In this section, we describe the steps and strategies used by our approach to generate such tests.
@@
Figure 2. Test code generated by Issue2Test for Django-12284 issue.
@@ 2.3. Phase 1: Issue Comprehension
IssueID: django-12284, Found files: 1. django/db/models/options.py 2. django/db/models/fields/__init__.py 3. django/db/models/fields/mixins.py 4. django/db/models/base.py 5. django/db/models/query.py
@@ 21             $isAssertionFailureRelated\leftarrow\textbf{assertion\textunderscore match}(failingTests,G)$
31       if ($failureType=\text{passing}$) or (not $isAssertionFailureRelated$) then
@@ 3.1.2. LLM Empowering Issue2Test
For all our experiments, we use OpenAI’s GPT-4o-mini (version 2024-07-18). GPT-4o-mini is smaller than GPT-4o but faster, offering a 128K context window. At the time of writing the paper, the used model costs \$0.15 per million input tokens and \$0.60 per million output tokens.
@@ 3.2.1. Issue2Test’s effectiveness
Table 1 presents the effectiveness of Issue2Test in generating F→P test cases per project. The results are reported separately for verified issues and all issues in the SWT-bench-lite dataset. Across all projects, Issue2Test achieves an overall F→P rate of 30.4%, successfully generating F→P tests for 84 out of 276 issues. On the verified subset, Issue2Test reproduces 30 out of 90 cases.
@@ 3.2.2. Comparison with Baselines
Issue2Test 84 30.4%
@@
Table 2 compares Issue2Test against the baselines (Section 3.1.4) using the same 276 instances from SWT-bench-lite. Among the baselines, ZeroShot achieves the lowest F→P rate, being successful for only 5.8% of the issues. AutoCodeRover and Libro perform better, achieving 9.1% and 15.2%, respectively. ZeroShotPlus improves upon ZeroShot by leveraging a custom diff format designed to be more robust for LLM-generated patches. It is worth noting that while AutoCodeRover and SWE-Agent are primarily designed for resolving GitHub issues through prompt modification, they are still capable of generating a moderate number of patch-validating tests. SWE-Agent outperforms these techniques with a 16.7% F→P rate, further improving to 19.2% with SWE-Agent+, which incorporates execution-based filtering. The improvement in SWE-Agent+ comes from the incorporation of execution-based filtering, where generated tests are executed on the pre-patch code, and those that do not fail as expected are discarded. However, no explicit refinement is performed to correct failing tests that do not fully capture the reported issue, limiting its ability to handle assertion mismatches or refine test logic for more precise issue reproduction. Auto-TDD achieved the highest F→P rate among prior methods, generating F→P tests for 21.7% of issues. Issue2Test outperforms Auto-TDD, achieving 30.4% and resolving 84 issues. Our approach differs from these techniques as we design a careful workflow that not only validates test behavior but also iteratively refines failing test cases. This enables Issue2Test to address assertion mismatches, resolve runtime failures, and improve test relevance, leading to a higher fail-to-pass (F→P) rate.
@@
Figure 16 presents a Venn diagram comparing the F→P tests generated by Issue2Test and four baseline techniques, including Libro, AutoCodeRover, SWE-Agent+, and Auto-TDD. The diagram shows both unique and overlapping F→P tests among these methods. Issue2Test demonstrates the highest number of unique tests, with 28 unique F→P tests not generated by any other approach. This indicates that Issue2Test can generate a significant number of tests that are not covered by existing methods. In contrast, Auto-TDD, SWE-Agent+, AutoCodeRover, and Libro each produce fewer unique test cases, ranging from 5 to 6. Overall, the results demonstrate that Issue2Test not only complements prior techniques, but also is the only technique to successfully handle a larger number of cases.
@@ 3.2.4. Example of Failure Cases
Figure 16. Venn diagram of F→P tests generated by different approaches.
@@ 3.3. Influence of Different Components of Issue2Test (RQ2)
Assertion matching exhibits the highest proportion of successful test cases relative to unsuccessful ones, indicating that when Issue2Test correctly identifies an issue-related assertion, it reliably generates a valid test. However, when the assertion does not align with the issue, additional refinement is required.
Error categorization plays a crucial role in distinguishing relevant failures from unrelated ones, with 13.5% of test cases successfully generated at this stage. The relatively low proportion of unsuccessful cases suggests that accurate failure classification significantly contributes to effective test generation.
Runtime and compilation error handling presents the most significant challenge, with 27.9% of test cases failing at this stage. This suggests that resolving test cases that crash or fail to compile is a major bottleneck, limiting the overall effectiveness of Issue2Test.
Test refinement, designed to improve failing test cases, results in a lower proportion of successful test cases (2.7%) compared to unsuccessful ones (4.8%), suggesting that refinement strategies could be further improved to better align test modifications with the reported issue.
@@ 5. Related Work
LLM-Based Test Generation. LLMs have been applied to test generation (Feng and Chen, 2024; Kang et al., 2023; Schäfer et al., 2023a, b; Nashid et al., 2023), with efforts to enhance quality by overcoming coverage plateaus (Lemieux et al., 2023), incorporating code-aware prompting (Ryan et al., 2024), and leveraging coverage information (Pizzorno and Berger, 2024). LLMs have also been explored for augmenting existing tests (Alshahwan et al., 2024) refines human-written test suites to improve coverage. Our work differs by focusing not on direct code generation but on synthesizing test cases specifically for reproducing issues raised on GitHub and revealing the buggy behavior related to the issue.
LLMs for GitHub Issue Resolution. LLMs are increasingly being leveraged to automate GitHub issue resolution by generating patches (Tao et al., 2025). Recent research has explored LLM-based agents that analyze issue reports, suggest code modifications, and improve software maintenance workflows. A key area of focus is automated program repair, where agents such as RepairAgent (Bouzenia et al., 2024) apply LLMs to detect and fix defects in source code. Other approaches, including SWE-Agent (Yang et al., 2024), MarsCode Agent (Liu et al., 2024), Magis (Tao et al., 2025), and AutoCodeRover (Zhang et al., 2024), tackle broader issue resolution tasks such as bug fixes, feature additions, and code enhancements. Our work, which automatically creates reproducing test cases for said issues, could help these approaches by: i) augmenting existing test suites and thus receiving more feedback about the behavior of code. ii) making the repair approaches more practical and realistic by using automatically generated test cases instead of developer-written tests, often, added after fixing the bug.

### v2

@@ Abstract.
Automated tools for solving GitHub issues are receiving significant attention by both researchers and practitioners, e.g., in the form of foundation models and LLM-based agents prompted with issues. A crucial step toward successfully solving an issue is creating a test case that accurately reproduces the issue. Such a test case can guide the search for an appropriate patch and help validate whether the patch matches the issue’s intent. However, existing techniques for issue reproduction show only moderate success. This paper presents Issue2Test, an LLM-based technique for automatically generating a reproducing test case for a given issue report. Unlike automated regression test generators, which aim at creating passing tests, our approach aims at a test that fails, and that fails specifically for the reason described in the issue. To this end, Issue2Test performs three steps: (1) understand the issue and gather context (e.g., related files and project-specific guidelines) relevant for reproducing it; (2) generate a candidate test case; and (3) iteratively refine the test case based on compilation and runtime feedback until it fails and the failure aligns with the problem described in the issue. We evaluate Issue2Test on the SWT-bench-lite dataset, where it successfully reproduces 32.9% of the issues, achieving a 16.3% relative improvement over the best existing technique. Our evaluation also shows that Issue2Test reproduces 20 issues that four prior techniques fail to address, contributing a total of 60.4% of all issues reproduced by these tools. We envision our approach to contribute to enhancing the overall progress in the important task of automatically solving GitHub issues.
@@ 1. Introduction
Issue reports, e.g., on GitHub, are commonly used to describe bugs, missing features, and other ways to improve a software project. Because addressing issues takes significant developer effort, recent work invests heavily in automated issue solving. Motivated by benchmarks of issues in popular open-source projects, such as SWE-bench (Jimenez et al., 2024), foundation models and Large Language Model (LLM)-based software engineering agents compete in their ability to successfully solve issues (Zhang et al., 2024; Roychoudhury et al., 2025; Rondon et al., 2025; Yang et al., 2024; Wang et al., 2025; Chen et al., 2024b; Tao et al., 2025; Xia et al., 2025; Pezzè et al., 2025). In addition to approaches that directly work on issues, automated program repair (Le Goues et al., 2019) tries to address bugs that manifest through a failing test. Various techniques have been explored for automated bug repair (Fan et al., 2023; Xia and Zhang, 2022; Bouzenia et al., 2025).
A crucial prerequisite for successfully solving an issue or repairing a bug is a test case that reproduces the problem. Unfortunately, issue reports rarely include executable tests that reproduce the issue, making it challenging for developers and automated techniques to validate potential patches. Specifically in open-source projects, contributors often lack the expertise and familiarity to construct meaningful test cases alongside their reports. Most work on automated program repair repair assumes reproducing test cases to be given. In practice, however, developers often write test cases only after a problem was fixed (Cheng et al., 2025). Automated issue solving techniques sometimes include a step for reproducing the issue (Ruan et al., 2025), but do not explicitly focus on this important part.
One approach for obtaining an issue-reproducing test case could be automated test generation. Both traditional approaches (Fraser and Arcuri, 2011; Pacheco et al., 2007) and neural methods (Watson et al., 2020; Dinella et al., 2022) attempt to reduce the time spent on manually writing tests. However, earlier techniques often fail to generate human-readable and maintainable test cases that align with project conventions and effectively validate intended program behavior (Daka et al., 2017). The emergence of LLMs has significantly enhanced automated test generation (Ziegler et al., 2022; Vaithilingam et al., 2022; Moradi Dakhel et al., 2023). LLMs can synthesize tests with human-like readability and writing style, reflecting patterns learned from extensive training on large code repositories (Ziegler et al., 2022). This capability has prompted the use of LLMs in test code generation (Bareiß et al., 2022; Nashid et al., 2023; Lemieux et al., 2023; Siddiq et al., 2023; Schäfer et al., 2023b; Chen et al., 2024a; Wang et al., 2024). While useful for creating regression tests, none of these approaches addresses the problem of creating tests that reproduce a specific issue.
@@
We evaluate Issue2Test on all 276 instances of SWT-bench-lite, a subset of SWE-bench that has been extensively used for evaluating prior work (Mündler et al., 2024; Ahmed et al., 2024). The results show that our approach: i) outperforms all the existing baselines by generating tests for 91 instances (13 more than the best baseline) with the best performing LLM, ii) successfully generates tests for 20 unique issues missed by any prior work and contributes 60.4% of the union of reproduced issues by any technique, iii) imposes moderate costs of 5.21 cents per issue, which is orders of magnitude cheaper than what a human software developer would charge for performing the same task.
@@ 2.1. Running Example
Title Cannot override get_FOO_display() in Django 2.2+ Description I cannot override the get_FIELD_display function on models since version $2.$ It works in version 2.1. Example: 1class FooBar(models.Model): 2 foo_bar = models.CharField(_("foo"), choices=[(1, ’foo’), (2, ’bar’)]) 3 4 def __str__(self): 5 return self.get_foo_bar_display() # This returns ’foo’ or ’bar’ in 2.2, but ’something’ in 2.1 6 7 def get_foo_bar_display(self): 8 return "something" What I expect is that I should be able to override this function.
@@
Figure 2 presents a test case generated by Issue2Test for the issue described in Figure 1, successfully reproducing the reported failure. Automatically generating such reproducing tests is practically valuable: it helps developers understand the issue more clearly, provides a concrete failure scenario to aid debugging, and enables patch validation in the absence of existing tests. For AI-based tools and agents, the generated test serves as an executable specification of the issue, potentially supporting automated repair and patch validation workflows.
While the running example involves a Django regression, many GitHub issues report failures where the underlying cause must be carefully validated. In such cases, prior approaches often produce tests that fail due to unrelated reasons, such as misconfigured environments or irrelevant assertions, without checking whether the failure corresponds to the reported issue. This underscores the need for verifying that generated test failures are semantically aligned with the issue description.
In this section, we describe the steps and strategies used by our approach to generate such tests.
@@
Figure 2. Test generated by Issue2Test for Django-12284.
@@ 2.3. Phase 1: Issue Comprehension
IssueID: django-12284 Found files: 1. django/db/models/options.py 2. django/db/models/fields/__init__.py 3. django/db/models/fields/mixins.py 4. django/db/models/base.py 5. django/db/models/query.py
@@ 21     $isAssertionFailureRelated\leftarrow\textbf{assertion\textunderscore match}(failingTests,G)$
31  if ($failureType=\text{passing}$) or (not isAssertionFailureRelatedisAssertionFailureRelated) then
@@ 3.1.2. LLM Empowering Issue2Test
For all our experiments, we employ GPT-4o-mini (version 2024-07-18), a cost-efficient proprietary model from OpenAI, as our primary model for evaluation. GPT-4o-mini is smaller than GPT-4o but faster, offering a 128K context window. At the time of writing the paper, the used model costs \$0.15 per million input tokens and \$0.60 per million output tokens.
To assess the impact of different models, we further evaluate Issue2Test with Meta Llama 3.3-70B (version meta.llama3-3-70b-instruct-v1:0 from AWS Bedrock) (Llama, 2025), an open-source model, and Claude 3.5 Sonnet (version 20241022) (Anthropic, 2025), a leading proprietary model from Anthropic.
@@ 3.1.4. Baselines
To compare Issue2Test with publicly available techniques, we include non-peer-reviewed entries from the SWT-Bench Lite leaderboard [^1] https://swtbench.com. This results in the inclusion of two variants of OpenHands (Wang et al., 2025), a generic-purpose coding agent. The first variant, OpenHands-Vanilla, uses the agent as released; the second variant, OpenHands-CI, augments the agent with a test‐code‐specific feedback extraction mechanism.
@@ 3.2.1. Issue2Test’s effectiveness
Table 1 presents the effectiveness of Issue2Test in generating F→P test cases per project using GPT-4o-mini. The results are reported separately for verified issues and all issues in the SWT-bench-lite dataset. Across all projects, Issue2Test achieves an overall F→P rate of 30.4%, successfully generating F→P tests for 84 out of 276 issues. On the verified subset, Issue2Test reproduces 30 out of 90 cases.
@@ 3.2.2. Comparison with Baselines
OpenHands-Vanilla (Claude 3-5 Sonnet) 63 22.8%
OpenHands-CI (Claude 3-5 Sonnet) 78 28.3%
Issue2Test (Llama 3.3) 49 17.8%
Issue2Test (GPT-4o-mini) 84 30.4%
Issue2Test (Claude 3-5 Sonnet) 91 32.9%
@@
Table 2 compares Issue2Test against the baselines (Section 3.1.4) using the same 276 instances from SWT-bench-lite. Among the baselines, ZeroShot achieves the lowest F→P rate, being successful for only 5.8% of the issues. AutoCodeRover and Libro perform better, achieving 9.1% and 15.2%, respectively. ZeroShotPlus improves upon ZeroShot by leveraging a custom diff format designed to be more robust for LLM-generated patches. It is worth noting that while AutoCodeRover and SWE-Agent are primarily designed for resolving GitHub issues through prompt modification, they are still capable of generating a moderate number of patch-validating tests. SWE-Agent outperforms these techniques with a 16.7% F→P rate, further improving to 19.2% with SWE-Agent+, which incorporates execution-based filtering. The improvement in SWE-Agent+ comes from the incorporation of execution-based filtering, where generated tests are executed on the pre-patch code, and those that do not fail as expected are discarded. However, no explicit refinement is performed to correct failing tests that do not fully capture the reported issue, limiting its ability to handle assertion mismatches or refine test logic for more precise issue reproduction. Auto-TDD achieved the highest F→P rate among prior methods, generating F→P tests for 21.7% of issues. Issue2Test outperforms Auto-TDD, achieving 30.4% and resolving 84 issues.
OpenHands-Vanilla achieves 22.8% F→P rate, demonstrating that a general-purpose agent without project-specific setup cannot extract test execution feedback during test generation. In contrast, OpenHands-CI is augmented with a continuous-integration pipeline: it installs all dependencies from the project manifest, injects the canonical test-suite invocation, and discards any diffs outside the existing tests/ directory. After each generated patch, the CI harness reruns the full suite and returns the failure trace. This structured feedback raises the F→P rate to 28.3 %, an improvement over the vanilla configuration. These results suggest that generic-purpose agents require task-specific adaptation for effective automated test generation.
[…]
Overall, Issue2Test differs from existing techniques by combining test validation with iterative refinement, resulting in more accurate and relevant reproducing tests. It analyzes whether failures, such as compilation errors, runtime exceptions, or assertion mismatches, align with the issue description, rather than treating any failure as sufficient. Diagnostic signals, including stack traces and assertion errors, are incorporated into a feedback loop that guides the LLM toward reproducing the issue. Furthermore, Issue2Test tailors test generation to project-specific conventions through meta-prompting, which is critical in frameworks where improper test scaffolding causes runtime errors, and it generates different test cases for mutually exclusive root-cause hypotheses, each reflecting a plausible explanation for the issue.
@@
Figure 16 presents a Venn diagram comparing the F→P tests generated by Issue2Test and four baseline techniques, including Libro, SWE-Agent+, Auto-TDD and OpenHands-CI. The diagram shows both unique and overlapping F→P tests among these methods. Issue2Test generates 20 unique F→P tests not generated by any other approach. This indicates that Issue2Test can generate a significant number of tests that are not covered by existing methods. In contrast, Auto-TDD, SWE-Agent+, AutoCodeRover, and Libro each produce fewer unique test cases, ranging from 5 to 6. Overall, the results demonstrate that Issue2Test not only complements prior techniques, but also is the only technique to successfully handle a larger number of cases.
@@ 3.2.4. Example of Failure Cases
Figure 16. Venn diagram of F→P tests across approaches.
@@ 3.3. Influence of Different Components of Issue2Test (RQ2)
Assertion matching exhibits the highest frequency in successful test cases relative to unsuccessful ones, indicating that when Issue2Test correctly identifies an issue-related assertion, it reliably generates a valid test. However, when the assertion does not align with the issue, additional refinement is required.
Error categorization plays a crucial role in distinguishing relevant failures from unrelated ones. It is invoked on average 13.5 times per issue, suggesting that repeated classification of test outcomes is central to the refinement loop. Accurate failure classification contributes significantly to effective test generation.
Runtime and compilation error handling presents a major challenge. It is invoked an average of 20.9 times per unsuccessful issue. This suggests that resolving test cases that crash or fail to compile is a significant bottleneck, limiting the overall effectiveness of Issue2Test.
Test refinement, designed to improve failing test cases, is invoked least often (2.7 times on average for successful cases and 4.8 for unsuccessful ones), suggesting that current refinement strategies may be less effective and could benefit from improved test transformation heuristics.
@@ 3.4. RQ3: Cost
4. Discussion
Quality of Generated Tests. Following studies on the quality of generated tests and their potential smells (Galindo-Gutierrez et al., 2023, 2025), we manually inspect a sample of successfully generated tests to check for smells and quality issues. First, we randomly sample, up to three successfully generated tests per project (some projects have fewer than three). This results in a total of 21 tests from 8 projects. Second, we manually go through the generated tests and their corresponding execution traces. For each generated test, we look for the categories of smells described in (Galindo-Gutierrez et al., 2023), namely: (1) Act–Assert Mismatch, (2) Redundant Code, (3) Failed Setup, (4) Accessors and Constants. We only find instances of redundant code in the inspected sample. Specifically, 5/21 tests had a duplicate setup, i.e., two or more test functions instantiate the same variables with identical values and identical subsequent changes to the variables. Furthermore, 2/21 tests had a duplicate test scenario, i.e, the same test scenario is implemented in two different test cases. The percentage of these two effects is similar to the values reported in prior work (Galindo-Gutierrez et al., 2023). On the positive side, Issue2Test does not show any smell symptoms from the other three categories. One reason for the duplication may be that our prompt instructs the LLM to distribute different test scenarios and covered branches over different functions, to isolate points of failure. Future work could investigate how to guide LLMs toward tests that avoid duplication and other test smells.
Reproduction Success by Issue Category. To assess whether Issue2Test depends on the kind of issue addressed, we classified all 276 SWT-bench-lite instances according to the SWE-bench taxonomy (Jimenez et al., 2024). The full set comprises 53 bug fixes (19.2%), 11 feature requests (4.0%), no regressions (0.0%), and 212 other issues (76.8 %), the latter dominated by maintenance-oriented tags such as “help wanted”. On the 84 issues for which Issue2Test generated reproducing tests, the distribution is 17 bugs (20.2 %), 4 features (4.8 %), 0 regressions, and 63 Others (75.0 %). These correspond to per-category success rates of 32.1% for bugs (17/53), 36.4% for feature requests (4/11), and 29.7% for other issues (63/212). The maximum gap across category is 6.7%, indicating that Issue2Test performs consistently on bug fixes, feature requests, and maintenance tasks. Improving reproduction accuracy for maintenance-style (“Other”) issues may require incorporating richer contextual information, such as more detailed task descriptions, to guide the synthesis of targeted and precise test cases.
Threats to Validity. One potential threat to internal validity is the evaluation dataset size. We evaluate Issue2Test on SWT-bench-lite, which consists of 276 issues from the larger SWT-bench dataset. While this benchmark has been widely used in prior studies, it may not fully capture the diversity of issue types encountered in practice (e.g, some projects in SWT-bench-lite have less than 10 issues). Another limitation of Issue2Test is its reliance on meta-prompting, which is particularly beneficial for popular open-source projects. However, while this approach enhances test generation accuracy, it raises concerns about Issue2Test’s ability to generalize to private repositories or lesser-known projects where LLMs lack prior exposure.
@@ 6. Related Work
LLM-Based Test Generation. LLMs have been applied to test generation (Feng and Chen, 2024; Kang et al., 2023; Schäfer et al., 2023a, b; Nashid et al., 2023), with efforts to enhance quality by overcoming coverage plateaus (Lemieux et al., 2023), incorporating code-aware prompting (Ryan et al., 2024), and leveraging coverage information (Altmayer Pizzorno and Berger, 2025). LLMs have also been explored for augmenting existing tests (Alshahwan et al., 2024) refines human-written test suites to improve coverage. Our work differs by focusing not on direct code generation but on synthesizing test cases specifically for reproducing issues raised on GitHub and revealing the buggy behavior related to the issue.
LLMs for GitHub Issue Resolution. LLMs are increasingly being leveraged to automate GitHub issue resolution by generating patches (Tao et al., 2025). Recent research has explored LLM-based agents that analyze issue reports, suggest code modifications, and improve software maintenance workflows. A key area of focus is automated program repair, where agents such as RepairAgent (Bouzenia et al., 2025) apply LLMs to detect and fix defects in source code. Other approaches, including SWE-Agent (Yang et al., 2024), MarsCode Agent (Liu et al., 2024), Magis (Tao et al., 2025), and AutoCodeRover (Zhang et al., 2024), tackle broader issue resolution tasks such as bug fixes, feature additions, and code enhancements. Our work, which automatically creates reproducing test cases for said issues, could help these approaches by: i) augmenting existing test suites and thus receiving more feedback about the behavior of code. ii) making the repair approaches more practical and realistic by using automatically generated test cases instead of developer-written tests, often, added after fixing the bug.
Execution-Oriented Agents. ExecutionAgent (Bouzenia and Pradel, 2024) and Issue2Test both use LLMs to automate software testing tasks but differ significantly in goals and techniques. ExecutionAgent focuses on automating the setup and execution of existing test suites by inferring and running project-specific build, dependency, and test commands. It assumes tests already exist and prioritizes execution without human input. In contrast, Issue2Test generates new test cases to reproduce failures described in issue reports, operating under the assumption that no reproducing test exists. It takes issue descriptions and stack traces as input to synthesize failing test logic. This means, while ExecutionAgent focuses on command line actions such as installing dependencies, Issue2Test, in contrast, focuses on generating test code. Subsequently, meta-prompting in ExecutionAgent helps infer commands and actions tied to the repository structure and languages while Issue2Test uses it to reflect project-specific test idioms such as naming, fixtures, and assertions. Though the principle of meta-prompting is the same, Issue2Test adapts its own meta-prompting for the target task. Furthermore, Issue2Test uses a similarity-based approach to find relevant test and code files while ExecutionAgent just relies on prompting and meta-promting to find relevant files at the start.
In addition, ExecutionAgent uses runtime feedback to fix command execution errors (e.g., build failures), refining its inferred commands. Issue2Test uses runtime diagnostics such as stack traces and assertion errors to iteratively refine the generated tests. It also introduces root cause hypothesis branching to explore alternative failure explanations and generate diverse test cases. This failure reasoning and root cause analysis approach is absent in ExecutionAgent.
On a high-level, while both systems rely on LLMs and meta-prompting, they serve complementary purposes: ExecutionAgent automates test execution, while Issue2Test enables semantic test synthesis from natural language and runtime signals.